OpenStudy (anonymous):

how do you factor this equation 1000x^3 +216 ??

OpenStudy (anonymous):

x^3=216/1000=0.216 x=0.46

OpenStudy (raden):

1000x^3 +216 = 10^3x^3 + 6^3 = (10x)^3 + 6^3 now, use the formula : a^3 + b^3 = (a+b)(a^2 - ab + b^2)